OpenStudy (anonymous):

solve 1) log(x+5)-log(x+1)=log(3x) 2) log5(x-6)=1-log5(x-2)

OpenStudy (shamil98):

Do you remember your log rules?

OpenStudy (shamil98):

\[\log_b M - \log_b N = \log_b \frac{ M }{ N }\]

OpenStudy (shamil98):

based off this rule what can you rewrite the first problem as?

OpenStudy (anonymous):

log x+5/x+1 = 3xlog

OpenStudy (anonymous):

?

OpenStudy (shamil98):

yes but it is log 3x

OpenStudy (shamil98):

log = log_10 so \[\log_{10} \frac{ x+5 }{ x+1 } = \log_{10} 3x\]

OpenStudy (anonymous):

oh okayy

OpenStudy (anonymous):

let me try again

OpenStudy (shamil98):

can you solve it from there?

OpenStudy (anonymous):

I believe so :) thanks. what about the 2nd one?

OpenStudy (shamil98):

Move the - log to the other side and use the rule: \[\log_b M + \log_b N = \log_b (M*N)\]

OpenStudy (shamil98):

\[\huge \log_5(x-6)=1-\log_5(x-2)\] \[\huge \log_5 (x-6) + \log_5(x-2) = 1\]

OpenStudy (anonymous):

so then would it be x-6+x-2=5? and so on?

OpenStudy (shamil98):

(x-6)(x-2) = 5^1

OpenStudy (anonymous):

oh okay got it !

OpenStudy (anonymous):

thank you so much :)

OpenStudy (shamil98):

glad to be of help :)